OpenStudy (anonymous):

Andy is building a model of a square pyramid for a class project. The side length of the square base is 11 inches and the slant height of the pyramid is 15 inches. What is the surface area of the model pyramid?

OpenStudy (austinl):

You have two distinct shapes. A triangle and a square. \[A_{triangle} = \frac{1}{2}\times b\times h\] \[A_{square} = l\times w\] You have 4 triangles with: b = 11 h = 15 Square l = 11 w = 11

OpenStudy (austinl):

Does this make sense?

OpenStudy (anonymous):

yeah thank you sooooo much

OpenStudy (austinl):

No problem. Glad I could help!